# Quantum-Enhanced Neural Exchange-Correlation Functionals

### v2

@@ II.3.3 Feature Maps for quantum models
$U_{f}(\bm{n})=\bigotimes e^{-\frac{i}{2}\varphi(n_{k})\hat{G}_{k}},$ (15)
where $\hat{G}$ are generators (possibly multi-qubit operators), and $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ is a data transformation function. We explore three established strategies for embedding data in quantum circuits. The choice of those strategies can be encouraged by viewing the feature maps selected for this work as Fourier series, which allows for a deeper understanding of their structure and behavior. According to Schuld et al.Schuld et al. (2021) such feature maps admits a Fourier series representation with a potentially exponential number of modes, more precisely $2^{N_{q}}$. This means that QNNs have an expressive power that grows exponentially with the number of qubits. As a result, there is a potential quantum advantage in representing XC functionals using these QNNs, which cannot be reliably approximated classically for large $N_{q}$andman et al. (2022)
@@
$U^{(1)}_{f,\mathrm{pr}}(n_{k})=\bigotimes e^{-\frac{i}{2}n_{k}\hat{Y}_{l}}.$ (16)
@@
$1$-local Chebyshev Feature Map
Chebyshev feature map utilizes an efficient polynomial basisSchuld et al. (2021):
$U^{(1)}_{f,\mathrm{ch}}(\bm{n})=\bigotimes e^{-k\cdot\mathrm{arccos}(n_{k})\hat{Y}_{k}}.$ (17)
[…]
@@
Finally, the amplitude feature map maps the density vector to a quantum state as $|\bm{n}\rangle=U^{(N_{\text{grid}})}_{f,\mathrm{a}}(\bm{n})|0\rangle$. However, this method presents two important drawbacks: the density is also normalized to a unit vector which suppresses the information of the number of electrons and, in general, an efficient implementation algorithm for the unitary $U^{(N_{\text{grid}})}_{f,\mathrm{a}}(\bm{n})$ does not exist for arbitrary vectors $\bm{n}$.
@@ II.3.4 Quantum-classical hybrid architectures
We propose to leverage advantages of classical NNs and combine them with quantum models. Notably, the number of elements $N_{\mathrm{grid}}$ in the electron density $\bm{n}$ scales linearly with the number of atoms. Dasgupta and Herbert (2017) Even for small diatomic molecules, there are hundreds of grid points for the smallest grid density available with the best integration grids Dasgupta and Herbert (2017). This poses an issue with the feature maps based on angle encoding since the number of qubits $N_{\mathrm{q}}$ required to embed the entire density $\bm{n}$ equals its size $N_{\mathrm{grid}}$, posing severe requirements for potential implementation on near-term quantum computers. Hence, we devise a model, Global QNN (GQNN), that is composed of two layers:
@@
with a product feature map and an alternate ladder ansatz. We first partition the input $\bm{n}$ in $N_{\mathrm{b}}=\lceil N_{\mathrm{grid}/}/L\rceil$ batches and then apply $f^{\mathrm{q}}_{\theta}$ to each batch. This results in $N_{\mathrm{b}}$ outputs that become the inputs of a classical function $f_{\mathrm{c}}$ defined as
@@ II.3.5 Quantum-inspired architectures
where the product feature map (Eq. (16)) is throughout all the layers. We do not employ Chebyshev embedding (Eq. (17)) due to detrimental normalization of the data required, which suppresses the information about the number of electrons in our architectures. Similarly to the hybrid architecture (GQNN), we consider quantum layers that consist of $L$-local QNNs, defined in Eq. (19), with a product feature map and alternate ladder ansatz defined in Appendix C. The input is partitioned in $\bm{n}$ in $N_{\mathrm{b}}=\lceil N_{\mathrm{grid}/}/L\rceil$ batches and then $f^{\mathrm{q}}_{\theta}$ is applied to each batch. This is repeated $N_{l}$ times until the output is a single scalar, which represents $E_{\mathrm{XC}}$. If not (i.e., odd number of inputs), the output of quantum layers is combined with an MLP as in Eq. (20) to output $E_{\mathrm{XC}}$.
@@ III.1 Implementation within differentiable chemistry frameworks
Li et al. proposed an important method to facilitate the generalization of neural XC functionals: using the KS iterative procedure as a regularizer. Li et al. (2021) The novel idea is to compute the energy loss at every KS iteration and the electron density loss at the last KS iteration, and then to differentiate through $N_{\mathrm{KS}}$ iterations (solutions of Eq. (1)) which would enable the model to converge faster to the reference, i.e., in fewer KS iterations, while regularizing its parameters. To allow this feature, a fully differentiable chemistry code in 1D in JAX Bradbury et al. (2018) was developed Li et al. (2021) with the aim of demonstrating the regularization capacity of the framework. Li et al. (2021) Since it reduces the number of dimensions (to 1D) and hence simplifies the problem, this is an ideal setup for investigation of the capacity of QNN-based models to represent XC functionals. We adopt this exact framework with minimal changes to hyperparameters for all our preliminary results in 1D (i.e., only linear molecules).
However, realistic chemistry simulations require a full 3D treatment of molecular systems. To address this demand, Kasim et al. proposed a 3D differentiable chemistry framework implemented entirely in PyTorch Paszke et al. (2017), which enabled accurate training of functionals for various diatomic systems. Kasim and Vinko (2021) However, both specialised frameworks of Li et al. and Kasim et al. are not used by computational chemists on a regular basis. Therefore, in this work, one of our goals is to contribute to an existing effort that focuses on making a popular chemistry framework, PySCF Sun et al. (2018), differentiable. This differentiable framework is known as PySCF with Automatic Differentiation (PySCFAD) Zhang and Chan (2022). It is similar to Li’s 1D framework, as it is also based on JAX, but offers the capability for an accurate treatment of chemical systems in 3D, making it suitable for a wide range of chemical simulations. Working within the scope of the PySCFAD framework, we replaced the calls to the standard XC functionals in differentiable SCF calculations used in KS-DFT (see Sec. II) by (quantum) NNs. This entails that we introduce necessary modifications to the already well-tested KS-DFT implementation.
@@
In addition, standard RKS-DFT implementations fail to consistently represent the ground state density, preventing the training of accurate XC models. They assume that the ground state density $n(\mathbf{r})$ can be accurately represented by a single determinant wave function of a non-interacting KS system with binary orbital occupation (0 or 1). However, the exact density of an interacting system, achievable theoretically via Full Configuration Interaction (FCI), allows for fractional occupation numbers. Helgaker et al. (2013) To address this limitation, especially evident in the challenging strongly-correlated regime, we adopt the algorithm of Chai. Chai (2012) Other approaches can be employed with an increased computational cost. Cances et al. (2003) Our choice was dictated by the simplicity of its integration in existing codes, with detailed proofs given in ref. Chai, 2012 and our implementation in Appendix A.
@@
    For 1D cases, we utilized the KS-DFT implementation by Li et al. Li et al. (2021)
@@
    For 3D cases, we employed the differentiable KS-DFT implementation from PySCFAD Zhang and Chan (2022), to which we added the FO method developed by Chai Chai (2012).
The 1D framework simplifies the training of neural functionals by limiting spatial degrees of freedom, although it restricts the study to linear systems. This setup provides an ideal environment to explore quantum models in a simplified context. Conversely, the 3D framework permits the evaluation of similar systems within a comprehensive representation, allowing for a realistic analysis. See Figs. 2 and 3(a) for an overview of the complete method.
Figure 2: Differentiable chemistry framework for training quantum-enhanced XC functionals. (1) The molecular dataset consists of reference energies $E_{\mathrm{ref}}$ and densities $\mathbf{n}_{\mathrm{ref}}$ generated by an accurate ab initio method. (2) A differentiable (Q)NN $f_{\theta}$ outputs the XC energy. The corresponding potential is obtained via (automatic) differentiation. The Fock equations in KS-DFT are iteratively solved to obtain new orbitals until self-consistency. Each KS iteration contributes to the loss $\mathcal{L}$, resulting in regularization. Li et al. (2021) Trainable parameters $\theta$ are updated by a classical optimizer. (3) The converged parameters $\theta^{*}$ define the trained functional $f_{\theta^{*}}$, which is subsequently used to evaluate the properties of any molecule in stage (4). Brown arrows in (2) indicate the gradient flow through the procedure. CPU and QPU denote classical (i.e., central and graphical) and quantum processing units.
@@ III.2 Training neural XC functionals
In this work, we adopt the standard supervised ML approach. We generate the training (reference) dataset $D=\{(E^{i}_{\mathrm{ref}},n^{i}_{\mathrm{ref}})\}_{i=1}^{N_{\mathrm{mol}}}$ using an accurate ab initio method for $N_{\mathrm{mol}}$ different molecules (also including different geometries of the same molecules). For the 1D implementation, we use the original training and test datasets by Li et al. Li et al. (2021)  composed of hydrogen atom-based systems, e.g., H2, H2H2 and H4. The data is obtained with DMRG simulations, where $N_{\mathrm{grid}}=513$ is fixed for all the systems. For the 3D implementation, we use the PySCF software package and the coupled cluster singles and doubles (CCSD) Bartlett and Musiał (2007) method to generate our data for H2, which is equivalent to FCI (an exact solution within a basis set) in this case. When more than two electrons are present, we suggest using the singles and doubles coupled cluster method with perturbative triples (CCSD[T]), which is viewed as the “gold standard” in computational chemistry. Bartlett and Musiał (2007) A Gaussian-type orbital basis set, 6-31G, is used in all calculations. For an integration grid, we use SG-3 with a Becke scheme to combine multi-atomic grids, such that for a given molecule at all geometries we obtain the same number of grid points. Dasgupta and Herbert (2017) The grid density is set to minimum to reduce the number of neural XC functional inputs, which simplifies the complexity of the problem. This entails $N_{\mathrm{grid}}=1192$ for H2 and $N_{\mathrm{grid}}=2384$ for planar H4 in the 3D setting. Given that we train models that allow only for a fixed number of inputs ($N_{\mathrm{grid}}$), grids have to be adapted (e.g., by pruning) to have the same number of grid points for all molecules in the dataset. Dasgupta and Herbert (2017) However, this is only a limitation imposed by simple prototypical (quantum) models that are investigated in this work. See Appendix B for additional clarifications about integration grids.
@@
where the expectation is taken on $N_{\mathrm{mol}}$ molecules of the training set, $\eta_{k}$ weights the contribution of every energy at each KS iteration and $w_{j}$ represents the integration grid weight. We use the weighting scheme introduced by Li et al. Li et al. (2021): $\eta_{k}=0.9^{N_{\mathrm{KS}}-k}H(k-10)$, where $H$ is the Heaviside function to tune the contribution of every loss term. This composite loss function includes two primary components: the density loss $\mathcal{L}_{n}$ and the energy loss $\mathcal{L}_{E}$. These components are essential for guiding the neural network to accurately produce electron densities and total energies, respectively. Specifically, the density loss assesses the accuracy of the Kohn-Sham electron density produced by the DFT pipeline compared to the reference electron density. Meanwhile, energy loss measures the precision of the Kohn-Sham energies generated by the neural network against the reference energies at each iteration $k$. The energy loss is weighted by $\eta_{k}$ to emphasize the importance of later iterations.
All derivatives of the loss function are computed via JAX automatic differentiation in this work. On quantum hardware, exact differentiation schemes such as parameter-shift rule Schuld et al. (2019) or its generalized version Kyriienko and Elfving (2021) can be used (see Appendix D). The loss function is then minimized to obtain the optimal model parameters
@@
that define the optimal XC correlation functional. In the following scenario, we outline the use of a gradient-based method for training, although it should be noted that our approach is also compatible with gradient-free techniques, which would, however, prevent regularization as in ref. Li et al., 2021.
At each converged SCF calculation, the gradient of loss $\mathcal{L}$ is performed via the automatic differentiation (or potentially parameter-shift rule on a quantum device). The gradient is then provided to the classical optimizer (Adam Kingma and Ba (2014) or L-BFGS-B Zhu et al. (1997)). The learning rate for the Adam optimizer can be decreased from $10^{-2}$ to $10^{-4}$ as the training progresses to fine-tune the results. Otherwise, L-BFGS-B and Adam are used with the standard hyperparameters.

### v3

@@ II.3.3 Feature maps for quantum models
$U_{f}(\mathbf{n})=\bigotimes_{k=1}^{N_{q}}e^{-\frac{i}{2}\varphi(n_{k})\hat{G}_{k}},$ (15)
where $\hat{G}$ are generators (possibly multi-qubit operators), and $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ is a data transformation function. We explore three established strategies for embedding data in quantum circuits. The choice of those strategies can be encouraged by viewing the feature maps selected for this work as Fourier series, which allows for a deeper understanding of their structure and behavior. According to Schuld et al.schuld2021effect such feature maps admit a Fourier series representation with a (potentially) exponential number of modes in the number of encoding gates or layers. This means that QNNs have an expressive power that may grow exponentially. It remains an open question whether the greater expressivity of QNNs provides a tangible advantage in representing XC functionals, even when such models cannot be reliably approximated classically for large $N_{q}$ landman2022classically. However, despite decades of research towards a universal XC functional, no single functional has proven accurate across all molecular systems. In this context, such new model types could offer an intriguing alternative.
@@
$U^{(1)}_{f,\mathrm{pr}}(n_{k})=\bigotimes_{l=1}^{N_{q}}e^{-\frac{i}{2}n_{k}\hat{Y}_{l}}.$ (16)
@@
11-local Chebyshev feature map
Chebyshev feature map utilizes a polynomial basisschuld2021effect:
$U^{(1)}_{f,\mathrm{ch}}(\mathbf{n})=\bigotimes_{k=1}^{N_{q}}e^{-i\cdot k\cdot\mathrm{arccos}(n_{k})\hat{Y}_{k}}.$ (17)
However, normalization of the density to a unit vecteor, i.e. $𝐧\mathbf{n}\leftarrow\mathbf$$|\mathbf{n}$$|$, is necessary. It suppresses electron count information and thus renders this approach unsuitable for density-based models.
@@
Finally, the amplitude feature map maps the density vector to a quantum state as $|\mathbf{n}\rangle=U^{(N_{\text{grid}})}_{f,\mathrm{a}}(\mathbf{n})|0\rangle$. However, this method presents two important drawbacks: the density is also normalized to a unit vector which suppresses the information of the number of electrons and, in general, an efficient implementation algorithm for the unitary $U^{(N_{\text{grid}})}_{f,\mathrm{a}}(\mathbf{n})$ does not exist for arbitrary vectors $\mathbf{n}$ nakaji2022approximate.
@@ II.3.4 Quantum-classical hybrid architectures
We propose to leverage advantages of classical NNs and combine them with quantum models. Notably, the number of elements $N_{\mathrm{grid}}$ in the electron density $\mathbf{n}$ scales linearly with the number of atoms. dasgupta2017standard Even for small diatomic molecules, there are hundreds of grid points for the smallest grid density available with the best integration grids dasgupta2017standard. This poses an issue with the feature maps based on angle encoding since the number of qubits $N_{\mathrm{q}}$ required to embed the entire density $\mathbf{n}$ equals its size $N_{\mathrm{grid}}$, posing severe requirements for potential implementation on near-term quantum computers. Hence, we devise a model, Global QNN (GQNN), that is composed of two layers:
@@
with a product feature map and an alternate ladder ansatz. We first partition the input $\mathbf{n}$ in $N_{\mathrm{b}}=\lceil N_{\mathrm{grid}/}/L\rceil$ batches and then apply $f^{\mathrm{q}}_{\theta}$ to each batch. This results in $N_{\mathrm{b}}$ outputs that become the inputs of a classical function $f_{\mathrm{c}}$ defined as
@@ II.3.5 Quantum-inspired architectures
where the product feature map (Eq. (16)) is throughout all the layers. We do not employ Chebyshev embedding (Eq. (17)) due to detrimental normalization of the data required, which suppresses the information about the number of electrons in our architectures. Similarly to the hybrid architecture (GQNN), we consider quantum layers that consist of $L$-local QNNs, defined in Eq. (19), with a product feature map and alternate ladder ansatz defined in Appendix C. The input is partitioned in $\mathbf{n}$ in $N_{\mathrm{b}}=\lceil N_{\mathrm{grid}/}/L\rceil$ batches and then $f^{\mathrm{q}}_{\theta}$ is applied to each batch. This is repeated $N_{l}$ times until the output is a single scalar, which represents $E_{\mathrm{XC}}$. If not (i.e., odd number of inputs), the output of quantum layers is combined with an MLP as in Eq. (20) to output $E_{\mathrm{XC}}$. Note that, although all models in this work are simulated, we classify this subset (QiNN and QiCNN) as quantum-inspired because the amplitude encoding used in the Global QiNN (i.e., $|\mathbf{n}\rangle=\sum_{j=1}^{N_{q}}\mathbf{n}_{j}|j\rangle$, where $\mathbf{n}$ is assumed to be normalized) typically requires an exponential number of gates zhang2021low; nakaji2022approximate, rendering it impractical on current NISQ devices. The QiCNN model is implemented as a sequence of QNNs in which the inputs to each layer are the expectation values from the previous layer, making the model more susceptible to error accumulation, unsuited for current noisy architectures.
@@ III.1 Implementation within differentiable chemistry frameworks
Li et al. proposed an important method to facilitate the generalization of neural XC functionals: using the KS iterative procedure as a regularizer. ksr The novel idea is to compute the energy loss at every KS iteration and the electron density loss at the last KS iteration, and then to differentiate through $N_{\mathrm{KS}}$ iterations (solutions of Eq. (1)) which would enable the model to converge faster to the reference, i.e., in fewer KS iterations, while regularizing its parameters. To allow this feature, a fully differentiable chemistry code in 1D in JAX jax2018github was developed ksr with the aim of demonstrating the regularization capacity of the framework. ksr Since it reduces the number of dimensions (to 1D) and hence simplifies the problem, this is an ideal setup for investigation of the capacity of QNN-based models to represent XC functionals. We adopt this exact framework with minimal changes to hyperparameters for all our preliminary results in 1D (i.e., only linear molecules).
However, realistic chemistry simulations require a full 3D treatment of molecular systems. To address this demand, Kasim et al. proposed a 3D differentiable chemistry framework implemented entirely in PyTorch paszke2017automatic, which enabled accurate training of functionals for various diatomic systems. kasim2021learning However, both specialised frameworks of Li et al. and Kasim et al. are not used by computational chemists on a regular basis. Therefore, in this work, one of our goals is to contribute to an existing effort that focuses on making a popular chemistry framework, PySCF sun2018pyscf, differentiable. This differentiable framework is known as PySCF with Automatic Differentiation (PySCFAD) zhang2022differentiable. It is similar to Li’s 1D framework, as it is also based on JAX, but offers the capability for an accurate treatment of chemical systems in 3D, making it suitable for a wide range of chemical simulations. Working within the scope of the PySCFAD framework, we replaced the calls to the standard XC functionals in differentiable SCF calculations used in KS-DFT (see Sec. II) by (quantum) NNs. This entails that we introduce necessary modifications to the already well-tested KS-DFT implementation.
@@
In addition, standard RKS-DFT implementations fail to consistently represent the ground state density, preventing the training of accurate XC models. Although the exact interacting ground-state wave function cannot, in general, be represented by a single-determinant wave function of a non-interacting KS system with binary orbital occupations (0 or 1), this is not a problem for the density itself. Still, RKS-DFT can fail numerically for open-shell systems (typically strongly correlated systems near dissociation), because the usual expression for the density as a sum over occupied orbitals becomes ill-defined: KS valence orbitals may be (near-)degenerate, which can lead in practice to convergence issues in (R)KS-DFT. This, in turn, necessitates fractional occupation of the KS system, which can no longer be described by a single Slater determinant but instead must be treated as an ensemble. To address this limitation, especially evident in the challenging strongly-correlated regime, we adopt the algorithm of Chai. frac-occ Other approaches can be employed with an increased computational cost. cances2003quadratically Our choice was dictated by the simplicity of its integration in existing codes, with detailed proofs given in ref. frac-occ and our implementation in Appendix A.
@@
    For 1D cases, we utilized the KS-DFT implementation by Li et al. ksr
@@
    For 3D cases, we employed the differentiable KS-DFT implementation from PySCFAD zhang2022differentiable, to which we added the FO method developed by Chai frac-occ.
The 1D framework simplifies the training of neural functionals by limiting spatial degrees of freedom, although it restricts the study to linear systems. This setup provides an ideal environment to explore quantum models in a simplified context. Conversely, the 3D framework permits the evaluation of similar systems within a comprehensive representation, allowing for a realistic analysis. To provide a quick overview, Figs. 2 and 3(a) summarize our framework in the following steps:
  1. (1)
    Generate dataset. Electronic structure calculations are performed via the best available level of theory (e.g., Full Configuration Interaction (FCI) helgaker2013molecular, DMRG) within a basis set (e.g., 6-31G* or larger) to create a dataset consisting of a set of energies and densities $\{(E^{i}_{\mathrm{ref}},\mathbf{n}^{i}_{\mathrm{ref}})\}_{i=1}^{N_{\mathrm{mol}}}$, with $\mathbf{n}^{i}_{\mathrm{ref}}$ evaluated on a selected grid of size $N_{\mathrm{grid}}$, for $N_{\mathrm{mol}}$ molecular geometries.
  2. (2)
    Train models. At every epoch, $N_{\mathrm{KS}}$ iterations are performed, each calling a neural functional $f_{\theta}$ that outputs the XC energy $E_{\mathrm{XC}}$ via local (Eq. (4)) or global embeddings (Eq. (5)). The corresponding $v_{\mathrm{KS}}$ potential is obtained via (automatic) differentiation. Then, the Fock equations are solved to obtain new orbitals. This whole procedure continues until self-consistency. Each KS iteration contributes to the loss $\mathcal{L}$, resulting in regularization. ksr At each epoch, trainable parameters $\theta$ are updated by a classical optimizer.
  3. (3)
    Select the best model. The converged parameters $\theta^{*}$ define the trained functional $f_{\theta^{*}}$ and are selected at one out of $N_{\mathrm{epochs}}$ epochs that provide the lowest loss on the validation set (i.e., selected geometries that are not present in the train and test sets).
  4. (4)
    Predict properties. The selected XC functional $f_{\theta^{*}}$ can be used to evaluate molecular properties. The neural functional can be completely agnostic to the input size ($N_{\mathrm{grid}}$) by design, and hence used for any molecular system.
Our implementation is available on GitHub as part of the QEX software package qex2025pasqal. In the next section, we describe each step of the framework in detail.
Figure 2: Differentiable chemistry framework for training quantum-enhanced XC functionals. (1) The molecular dataset consists of reference energies $E_{\mathrm{ref}}$ and densities $\mathbf{n}_{\mathrm{ref}}$ generated by an accurate ab initio method. (2) A differentiable (Q)NN $f_{\theta}$ outputs the XC energy. The corresponding potential is obtained via (automatic) differentiation. The Fock equations in KS-DFT are iteratively solved to obtain new orbitals until self-consistency. Each KS iteration contributes to the loss $\mathcal{L}$, resulting in regularization. ksr Trainable parameters $\theta$ are updated by a classical optimizer. (3) The converged parameters $\theta^{*}$ define the trained functional $f_{\theta^{*}}$, which is subsequently used to evaluate the properties of any molecule in stage (4). Brown arrows in (2) indicate the gradient flow through the procedure. CPU and QPU denote classical (i.e., central and graphical) and quantum processing units.
@@ III.2 Training neural XC functionals
In this work, we adopt the standard supervised ML approach. We generate the training (reference) dataset $D=\{(E^{i}_{\mathrm{ref}},n^{i}_{\mathrm{ref}})\}_{i=1}^{N_{\mathrm{mol}}}$ using an accurate ab initio method for $N_{\mathrm{mol}}$ different molecules (also including different geometries of the same molecules). For the 1D implementation, we use the original training and test datasets by Li et al. ksr  composed of hydrogen atom-based systems, e.g., H2, H2H2 and H4. The data is obtained with DMRG simulations, where $N_{\mathrm{grid}}=513$ is fixed for all the systems. For the 3D implementation, we use the PySCF software package and the coupled cluster singles and doubles (CCSD) bartlett2007coupled method to generate our data for H2, which is equivalent to FCI (an exact solution within a basis set) in this case. When more than two electrons are present, we suggest using the singles and doubles coupled cluster method with perturbative triples (CCSD[T]), which is viewed as the “gold standard” in computational chemistry for weakly correlated systems. bartlett2007coupled In the presence of strong correlation, the aforementioned CC methods often fail to converge to the correct solution. bulik2015can Consequently, we employ FCI in challenging cases, such as the square H4 system, for which it remains feasible given the small system sizes considered in this work. A Gaussian-type orbital basis set, 6-31G, is used in all calculations. For an integration grid, we use SG-3 with a Becke scheme to combine multi-atomic grids, such that for a given molecule at all geometries we obtain the same number of grid points. dasgupta2017standard The grid density is set to minimum to reduce the number of neural XC functional inputs, which simplifies the complexity of the problem. This entails $N_{\mathrm{grid}}=1192$ for H2 and $N_{\mathrm{grid}}=2384$ for planar H4 in the 3D setting. Given that we train models that allow only for a fixed number of inputs ($N_{\mathrm{grid}}$), grids have to be adapted (e.g., by pruning) to have the same number of grid points for all molecules in the dataset. dasgupta2017standard However, this is only a limitation imposed by simple prototypical (quantum) models that are investigated in this work. See Appendix B for additional clarifications about integration grids.
@@
where the expectation is taken on $N_{\mathrm{mol}}$ molecules of the training set, $\eta_{k}$ weights the contribution of every energy at each KS iteration and $w_{j}$ represents the integration grid weight. We use the weighting scheme introduced by Li et al. ksr: $\eta_{k}=0.9^{N_{\mathrm{KS}}-k}H(k-10)$, where $H$ is the Heaviside function to tune the contribution of every loss term. This composite loss function includes two primary components: the density loss $\mathcal{L}_{n}$ and the energy loss $\mathcal{L}_{E}$. These components are essential for guiding the neural network to accurately produce electron densities and total energies, respectively. Specifically, the density loss assesses the accuracy of the Kohn-Sham electron density produced by the DFT pipeline compared to the reference electron density. Meanwhile, energy loss measures the precision of the Kohn-Sham energies generated by the neural network against the reference energies at each iteration $k$. The energy loss is weighted by $\eta_{k}$ to emphasize the importance of later iterations.
All derivatives of the loss function are computed via JAX automatic differentiation in this work. On quantum hardware, exact differentiation schemes such as parameter-shift rule schuld2019evaluating or its generalized version kyriienko2021generalized can be used (see Appendix D). The loss function is then minimized to obtain the optimal model parameters
@@
that define the optimal XC correlation functional. In the following scenario, we outline the use of a gradient-based method for training, although it should be noted that our approach is also compatible with gradient-free techniques, which would, however, prevent regularization as in ref. ksr.
At each converged SCF calculation, the gradient of loss $\mathcal{L}$ is performed via the automatic differentiation (or potentially parameter-shift rule on a quantum device). The gradient is then provided to the classical optimizer (Adam kingma2014adam or L-BFGS-B zhu1997algorithm). The learning rate for the Adam optimizer can be decreased from $10^{-2}$ to $10^{-4}$ as the training progresses to fine-tune the results. Otherwise, L-BFGS-B and Adam are used with the standard hyperparameters.